OpenStudy (anonymous):

Use rational exponets to write x (1/8) * y(1/5) * z(1/2) as single radical expression.

OpenStudy (anonymous):

yikes. like adding fractions, need a common denominator of 40 so \[\frac{1}{8}=\frac{5}{40}\] \[\frac{1}{5}=\frac{8}{40}\] \[\frac{1}{2}=\frac{20}{40}\] and your answer is \[\sqrt[40]{x^5y^8z^{20}}\]